Outdoor at 5°CDBT and 60%RH is passed through a heating coil which gives 11kW
energy to the air, then humidified by sat steam (at 110oC) with a rate of 0.33kg/min, find
the final air dry and wet bulb temperatures for unit mass rate.
1
Expert's answer
2020-07-27T07:23:17-0400

Here in the question , DBT="5^oC" , RH=60 %,

from the psychometric chart we find

H=15 J/g, Humidity ratio=0.03

humidified by steam at 110 oC and mass flow rate is 0.33 kg/min

Total energy supplied=11kJ/s

water removed in dryer= "\frac {0.33}{60}=0.055 kg/s"

from psychometric

Total energy supplied = 268-35=233

so final condition for air

DBT=35oC, WBT= 33 oC using psychometric